# Regularizing Extrapolation in Causal Inference

### v1

@@ 6 Conclusion
This work proposes a unified framework for regularizing extrapolation in causal inference by replacing hard non-negativity constraints with soft penalties on negative weights. Our approach reveals a fundamental “bias-bias-variance” tradeoff between distributional imbalance, model misspecification, and estimator variance, with theoretical error bounds that decompose extrapolation bias through a novel reflection perspective. Empirical studies on synthetic data and a real-world medication trial demonstrate that controlled extrapolation can outperform both fully constrained and unconstrained approaches, particularly in high-dimensional settings with poor positivity. Our approach focuses primarily on weighting-type estimators, leaving open questions about how our results extend to other estimator classes. Second, our theoretical guarantees rely on Hölder continuity assumptions for expected outcomes, which may not hold in all practical settings. Future research directions include extending the bias-bias-variance tradeoff analysis to a broader class of estimatorsand exploring weaker or alternative continuity assumptions that might better capture real-world outcome functions. The framework presented here represents an important step toward more nuanced approaches to positivity violations in causal inference, moving beyond binary perspectives on extrapolation toward a continuous spectrum of regularization strategies.
@@ Appendix A Implementational Details
We implement the methods and case studies in this paper using Python 3.10. We implemented our weight estimation framework using PyTorch (version 2.7.0) for efficient automatic differentiation and optimization. The optimization is performed using Adam optimizer with defaults learning rate of 0.01 for default of 5,000 epochs. By default the weights are normalized to sum to one after each optimization step – however, user can choose otherwise. The implementation includes comprehensive visualization tools, including Love plots for covariate balance assessment and 2D scatter plots with convex hull visualization for geometric interpretation. All random operations are seeded for reproducibility, and the code supports both single and multiple outcome variables. For the MOUD case study in Section D, we scale the pre-treatment data to ensure that the maximum value for each covariate is 1 and the minimum is 0; it is important to note that most covariates in this instance are discrete binary.
Appendix B Proofs
B.1 Proof of Proposition 3.1
@@ Appendix B Proofs
Appendix C Synthetic Data Study
In this section, we study the performance of our estimator using two synthetic studies, one using a linear data generative process (DGP) and the second one using a non-linear data generative process (DGP). In this simulation study, we specifically consider a case when the target point $\boldsymbol{x}$$^{\star}$ is outside the convex hull of the training points $\{{\boldsymbol{X}}_{1},\dots,{\boldsymbol{X}}_{n}\}$. For linear DGP, the outcome $Y$ is a linear function of $\boldsymbol{X}$} while for nonlinear DGP, the outcome is a quadratic function of $\boldsymbol{X}$}. In particular, the DGPs are as follows:
@@ Appendix C Synthetic Data Study
Here, we consider a sample of 10 training units and a target unit as shown in Figure 6. This scenario is particularly interesting because of the limited sample size compared to the problem’s dimensionality.
@@
For the linear data-generating process, our parametric assumption holds. We observe that increasing the regularization on extrapolation ($\gamma$) decreases the negative influence while increasing the balance RMSE, as shown in Figures 7(b) and (c) – consistent with theoretical expectations. As underlying DGP is linear, relying on parametric assumptions for extrapolation yields optimal estimates with the smallest estimation error corresponding to least level of regularization on extrapolation (see Figure 7(a)).
@@
Appendix D Generalizing Medication for Opioid Use Disorder Trial Evidence
Data Description.
The Starting Treatment With Agonist Replacement Therapies (START) trial, initiated in 2006, was a multi-center study comparing buprenorphine versus methadone in treating opioid use disorder [Saxon et al., 2013, Hser et al., 2014]. The trial enrolled 1,271 participants, who were randomized in a 2:1 ratio to receive either buprenorphine or methadone. Methadone was found to have higher rates of patient retention in treatment compared to buprenorphine (though buprenorphine in this trial was given in an unusual way to mimic methadone medication administration—requiring near daily clinic visits of participants) [Hser et al., 2014]. Our analysis focuses on the outcome of relapse to regular opioid use within 24 weeks of medication assignment, defined as non-study opioid use for four consecutive weeks or daily use for seven consecutive days. Data on opioid use were collected through urine drug screens and self-reports, with relapse assessment beginning 20 days post-randomization to account for residual drug presence during stabilization.
Parikh et al. [2025] identified that Latina women with a pre-treatment history of amphetamine and benzodiazepine use were underrepresented in the START trial relative to the target population, highlighting a practical violation of the positivity assumption. In this study, we estimate the average treatment effects (ATE) for this underrepresented subgroup using our proposed framework alongside standard linear regression and inverse probability weighting (IPW) estimators. We also evaluate the sensitivity of these estimates to parametric assumptions.
The target sample is drawn from the 2015–2017 Treatment Episode Dataset - Admissions (TEDS-A), which includes data on individuals entering publicly funded substance use treatment programs across 48 states (excluding Oregon and Georgia) and the District of Columbia. Our analysis focuses on Latina women with a pre-treatment history of amphetamine and benzodiazepine use.
Analysis.
Using linear regression, the estimated treatment effect for our subgroup of interest is 0.278-0.278, indicating that relapse rates are approximately 28 percentage points lower under methadone compared to buprenorphine. In contrast, the IPW estimate is 0.014-0.014, suggesting that both treatments are similarly effective. However, as shown in Figure 11, the negative influence of the linear regression-based estimate is 35% and 40% for $T=0$ and $T=1$, respectively, compared to 0% for IPW. Despite this, Figure 10 demonstrates that the IPW estimator achieves worse covariate balance than linear regression. These findings reveal that while linear regression relies heavily on parametric assumptions and negative weights, it may be biased if these assumptions are violated. Conversely, the IPW estimator avoids additional parametric assumptions but introduces bias due to poor covariate balance and violations of the positivity assumption.
Figure 9: Average Treatment Effects for the Target Sample from TEDS-A of Hispanic Females who have a history of Amphetamize and Benzodiazepine use. Each hue corresponds to a value of $\lambda$ and the x-axis corresponds to different values of $\gamma$ (on log scale). The dashed line represents the estimate using linear regression and the dotted represents the estimate using inverse probability weighting (IPW).
We apply our proposed framework to address these issues, which regularizes extrapolation to mitigate reliance on extreme weights. By varying $\gamma$ from 0.01 to 10, we examine how treatment effect estimates shift with increasing regularization of negative weights. Without regularization, the estimates converge with those from linear regression. However, as regularization intensifies, the estimates smoothly shift towards zero and occasionally change the sign from negative to positive for smaller values of $\lambda$. This sensitivity underscores the influence of assumptions on the point estimates. While increasing $\gamma$ reduces negative influence (Figure 11), it worsens covariate balance, as reflected in higher RMSE values (Figure 10). Thus, our framework highlights a trade-off between minimizing reliance on parametric assumptions and achieving optimal covariate balance.
Figure 10: Balance between the trial and the target samples measured as the root mean squared error (RMSE) for different values of $\gamma$ and $\lambda$ along with implied linear regression weights and inverse probability weights (IPW).
These findings suggest that applied researchers should interpret treatment effect estimates among under-represented subgroups with caution, given their sensitivity to modeling assumptions. As Parikh et al. [2025] emphasized, collecting more representative trial data is critical to credibly estimate treatment effects for this underrepresented subgroup in future medication for opioid use disorder studies.
Figure 11: Negative influence, defined as the contribution of negative weights in estimation, for different values of $\gamma$ and $\lambda$ along with implied linear regression weights and inverse probability weights (IPW).

### v2

@@ 5 Conclusion
This work proposes a framework for regularizing extrapolation in causal inference by replacing hard non-negativity constraints with soft penalties on negative weights. Our approach reveals a fundamental “bias-bias-variance” tradeoff between distributional imbalance, model misspecification, and estimator variance. We provide theoretical error bounds that decompose extrapolation bias through a novel reflection perspective, and demonstrate empirically on both synthetic data and a real-world medication trial that controlled extrapolation can outperform both fully constrained and unconstrained approaches, particularly in high-dimensional settings with poor positivity.
Limitations and Future Work. Our approach focuses primarily on weighting-type estimators and relies on Hölder continuity assumptions for expected outcomes and conditional ignorability, which may not hold in all practical settings. Future research should extend the bias-bias-variance tradeoff analysis to broader estimator classes, explore weaker continuity assumptions, and most importantly, operationalize sensitivity analysis for unmeasured confounding. While existing proposals for sensitivity analysis with balancing weights (Soriano et al., 2023) cannot be applied directly to our framework, adapting such methods represents a critical next step for practical implementation.
The framework presented here represents an important step toward more nuanced approaches to positivity violations in causal inference, moving beyond binary perspectives on extrapolation toward a continuous spectrum of regularization strategies that can be tailored to specific research contexts and risk tolerances.
@@ Appendix A Implementational Details
We implement the methods and case studies in this paper using Python 3.10. We implemented our weight estimation framework using PyTorch (version 2.7.0) for efficient automatic differentiation and optimization. The optimization is performed using Adam optimizer with defaults learning rate of 0.01 for default of 5,000 epochs. By default the weights are normalized to sum to one after each optimization step – however, user can choose otherwise. The implementation includes comprehensive visualization tools, including Love plots for covariate balance assessment and 2D scatter plots with convex hull visualization for geometric interpretation. All random operations are seeded for reproducibility, and the code supports both single and multiple outcome variables. For the MOUD case study in Section 4, we scale the pre-treatment data to ensure that the maximum value for each covariate is 1 and the minimum is 0; it is important to note that most covariates in this instance are discrete binary. Computational details: Our objective mimics a classic quadratic program (QP). QPs are well studied in the literature, and their run time complexity is known to be $O(n^{3.67}*logn)$ where $n$ is the number of training units [Kapoor and Vaidya, 1986]. Our new simulation study also demonstrates that our method is scalable to high-dimensional data.
Appendix B Connecting with Existing Estimators
Our proposed framework establishes meaningful connections with existing weighting and doubly robust estimators. When $\gamma=0$ and $p=2$, our estimator reduces to single ridge regression with hyperparameter $\lambda$. In this context, the parameter $\gamma$ can be interpreted as controlling the degree of extrapolation in the doubly robust (DR) estimator; see Ben-Michael et al. [2021b]. This relationship can be further understood through the lens of ridge-augmented balancing weights, which constitute a doubly robust estimator under certain regularity conditions [Bruns-Smith et al., 2023]. Building on the work of Knaus [2024], which demonstrated that many common causal inference estimators—including doubly robust methods like AIPW—can be expressed as weighting estimators, we can extend our framework to accommodate and regularize many existing approaches. Specifically, we can adapt our framework to work with implied weights from baseline estimators by modifying the original objective function as follows:
$\mathbf{w}^{*}\in\arg\min_{\mathbf{w}}\|\mathbf{w}-\mathbf{w}^{\prime}\|+\gamma\|\mathbbm{1}(w_{i}<0)|w_{i}|\left(\|{\boldsymbol{X}}_{i}\|^{\alpha}_{2}\right)\|$
where $\mathbf{w}^{\prime}$ represents the implied weights from a baseline estimator. This formulation allows us to regularize extrapolation in existing estimators such as AIPW and IPW, providing a principled approach to control their behavior in regions where overlap is limited.
Appendix C Proofs
C.1 Proof of Proposition 3.1
@@ Appendix C Proofs
C.3 Proof of proposition 3.3
Proof.
By Proposition 3.2, with probability at least $1-\delta$:
We bound $\|\hat{w}\|_{2}$ using the first-order optimality conditions. For each $i$:
$2\sum_{j=1}^{n}\hat{w}_{j}X_{j}^{T}X_{i}-2(x^{*})^{T}X_{i}+2\lambda\hat{w}_{i}+\gamma\|X_{i}\|^{\alpha}\xi_{i}=0$ (5)
where $\xi_{i}\in\partial(|w_{i}|\mathbf{1}(w_{i}<0))|_{w_{i}=\hat{w}_{i}}$. Multiplying by $\hat{w}_{i}$ and summing over all $i$:
$2\left\|\sum_{j=1}^{n}\hat{w}_{j}X_{j}\right\|_{2}^{2}-2(x^{*})^{T}\sum_{i=1}^{n}\hat{w}_{i}X_{i}+2\lambda\|\hat{w}\|_{2}^{2}-\gamma\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)\|X_{i}\|^{\alpha}=0$ (6)
Rearranging and applying Young’s inequality $(x^{*})^{T}\sum_{i}\hat{w}_{i}X_{i}\leq\frac{1}{2}\|x^{*}\|_{2}^{2}+\frac{1}{2}\|\sum_{i}\hat{w}_{i}X_{i}\|_{2}^{2}$:
$\lambda\|\hat{w}\|_{2}^{2}\leq\frac{1}{2}\|x^{*}\|_{2}^{2}+\frac{\gamma R^{\alpha}}{2}\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)$ (7)
By Cauchy-Schwarz, $\sum_{i=1}^{n}|\hat{w}_{i}|\mathbf{1}(\hat{w}_{i}<0)\leq\sqrt{n}\|\hat{w}\|_{2}$. This gives:
$\|\hat{w}\|_{2}^{2}-\frac{\gamma R^{\alpha}\sqrt{n}}{2\lambda}\|\hat{w}\|_{2}-\frac{\|x^{*}\|_{2}^{2}}{2\lambda}\leq 0$ (8)
Solving this quadratic inequality:
$\|\hat{w}\|_{2}\leq\frac{\gamma R^{\alpha}\sqrt{n}}{4\lambda}+\sqrt{\frac{\gamma^{2}R^{2\alpha}n}{16\lambda^{2}}+\frac{\|x^{*}\|_{2}^{2}}{2\lambda}}\leq\frac{\|x^{*}\|_{2}}{\sqrt{\lambda}}+\frac{\gamma R^{\alpha}\sqrt{n}}{\lambda}$ (9)
where we used $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ and simplified. Substituting into Proposition 3.2 completes the proof. ∎
Appendix D Synthetic Data Study
D.1 Challenge Study
In this section, we study the performance of our estimator via two data generative process (DGP), one using a linear DGP and the second one using a non-linear DGP. In this simulation study, we specifically consider a challenging case when the target point $\boldsymbol{x}$$^{\star}$ is outside the convex hull of the training points $\{{\boldsymbol{X}}_{1},\dots,{\boldsymbol{X}}_{n}\}$. For linear DGP, the outcome $Y$ is a linear function of $\boldsymbol{X}$} while for nonlinear DGP, the outcome is a quadratic function of $\boldsymbol{X}$}. The two DGPs are as follows:
@@ D.1 Challenge Study
To further make the setup more challenging, we consider a relatively small sample with 10 training units and a target unit as shown in Figure 4. This scenario is particularly interesting because of the limited sample size compared to the problem’s dimensionality: $n/p=5$.
@@
For the linear DGP, our parametric assumption holds. We observe that increasing the regularization on extrapolation ($\gamma$) decreases the negative influence while increasing the balance RMSE, as shown in Figures 5(b) and (c) – consistent with theoretical expectations. As underlying DGP is linear, relying on parametric assumptions for extrapolation yields optimal estimates with the smallest estimation error corresponding to least level of regularization on extrapolation (see Figure 5(a)).
@@
D.2 High-dimensional Large Scale Study
We now expand our analysis to a large scale high-dimensional data. In particular, we consider a scenario where $n/p$ is large. We generate synthetic data via Friedman DGP which is commonly used across machine learning and causal inference literature. We generate $n_{control}=1000$ control units and $n_{treat}=1000$ with $p=5000$ covariates. The DGP is Friedman DGP is defined as follows:
$Y(0)=10*\sin(\pi X_{0}X_{1})+20(X_{2}-0.5)^{2}+10X_{3}+5X_{4}+\epsilon,$
and $Y(1)=Y(0)$. In this example, we are interested in estimating average treatment effect on the treated – the truth is 0. Here, the control units are used as the training set estimate counterfactual for the treated units which act as estimation set here – $n/p=1000/5000=0.2$. We compare our approach with IPW, AIPW and Random Forest (RF). One main difference between our framework and AIPW or RF is that AIPW and RF both uses outcome information to learn the estimator while our framework (similar to IPW) is outcome data agnostic and only uses covariate information to ensure balance while also regularizing extrapolation. Figure 7 shows the mean squared error in estimating ATT and compares it with IPW, AIPW and RF. Our results shows that sweeping over the range of $\gamma$ (i.e. regularizing extrapolation) smoothly ‘interpolates’ between AIPW and IPW estimators (former allows extrapolation using linear regression outcome model while other only relies on interpolation.)
Figure 7: Results on synthetic data generated using High-dimensional Friedman’s DGP.